OpenStudy (anonymous):

If a= 2i +7j, find |a|.

OpenStudy (anonymous):

\[\sqrt{4+49}\]=\[\sqrt{53}\]

OpenStudy (anonymous):

I a I.... this called the norm of ( a ) I a I = ( 2^2 + 7^2 )^1/2 = (53)^1/2